# Curve-Aware Gaussian Splatting for 3D Parametric Curve Reconstruction

### v2

@@ 1 Introduction
Figure 1: Our method achieves clean and accurate parametric curve reconstruction through a one-stage optimization utilizing the proposed curve-aware Gaussians. It eliminates the limitations of existing two-stage approaches, where the accuracy of the stage-2 curve fitting is heavily dependent on the quality of the initial point cloud reconstruction in stage 1.
@@
Parametric curves, as fundamental geometric primitives, play a vital role in structured 3D representation and geometric understanding. Their ability to compress geometric information while maintaining CAD compatibility makes them indispensable for industrial applications. A critical challenge lies in reconstructing high-precision parametric curves from unstructured data, such as noised point clouds or multi-view images, as sensor imperfections and incomplete observations inevitably degrade reconstruction quality.
@@
Traditional methods for 3D curve reconstruction typically require clean 3D point clouds as input for subsequent fitting procedures. The recent NEF framework [35] first achieved 3D curve reconstruction directly from 2D images through a two-stage “edge point cloud reconstruction + parametric curve fitting” pipeline that combines neural radiance fields with RANSAC-based optimization. However, this approach accumulates errors from the disconnected two-stage optimization process, making it sensitive to the noise of the reconstructed point cloud. EMAP [12] improves upon NEF by enhancing edge localization through unbiased distance UDF modeling and elimination equation rendering. Though its “cluster-then-fitting” strategy reduces error propagation, the method remains fundamentally constrained by the point cloud quality given in the first stage. Notably, these approaches suffer from low computational efficiency due to independent neural optimizations at each stage. While EdgeGaussians [5] accelerates reconstruction using Gaussian splatting, it still fails to overcome the inherent limitations of a two-stage fashion pipeline.
@@
The two-stage paradigm introduces two fundamental limitations. First, the parametric fitting process is strongly coupled with observation noises/errors, propagated from 2D edge detection to 3D point clouds through multi-view triangulation. It results in redundant branches or breaks in final curves (as shown in Fig. 1). Secondly, greedy iterative fitting (e.g., RANSAC) produces local optima accumulation, causing curve redundancy to grow exponentially with scene complexity, which severely impacts downstream application efficiency. Eliminating the optimization gap between stages emerges as the key to robust reconstruction.
@@
While 2D edge maps are much easier to acquire, bypassing the explicit 3D edge point cloud reconstruction is a promising way to solve this problem. Direct one-stage curve optimization confronts two fundamental challenges: parametric representations inherently lack differentiable rendering capability, while neural rendering frameworks (NeRF/Gaussians) cannot preserve geometric continuity when approximating curves. We present a hybrid neural representation, CurveGaussians, that combines parametric curves with edge-oriented Gaussian components. By establishing bi-directional coupling, each Bézier curve segment precisely controls the spatial distribution of its associated Gaussian components. It enables differentiable rendering through splatting while significantly reducing the parameter count required for the optimization.
@@
Building upon this hybrid representation, this paper proposes a one-stage method for parametric curve reconstruction from multi-view edge maps. The core concept bypasses intermediate point cloud representations to directly perform differentiable optimization on the control points of Bézier curves, where each curve comprises multiple edge-oriented Gaussians. Specifically, each Gaussian component is bound to a specific edge segment of its corresponding curve, with positional and orientational attributes analytically derived from curve parameters. Rendering losses are backpropagated to update curve control points, enabling simultaneous exploitation of neural rendering’s photometric consistency constraints and parametric curves’ intrinsic geometric regularization. To further optimize curve topology during training for cleaner and more compact reconstructions, we propose a dynamic topology optimization strategy that adaptively refines edge structures through merging, splitting, and pruning operations. Additionally, we improve the optimization loss to prioritize the structural compactness of generated curves. Experimental results demonstrate that our one-stage method increases 14.5% in reconstruction accuracy, improves computational efficiency by 33%, and reduces generated curve counts by 70.5% compared with two-stage SOTA alternatives, attaining more accurate and compact reconstructions. In summary, the contributions include:
@@ 2 Related Works
Early methods [25, 23, 4] reconstruct 3D line segments using Structure-from-Motion (SfM), involving line detection [19, 20, 32], matching [1, 21], and triangulation [3, 2, 4, 15]. Other works address line mapping by integrating graph clustering [7] and planar constraints [29] into SLAM systems [14, 16, 24]. NEAT [33] optimizes 3D line junctions using VolSDF [34] but is limited to straight lines and textured objects. However, these methods are limited to straight lines, often producing fragmented segments for curves. Recent advances in 3D parametric curve reconstruction focus on two paradigms: point cloud-based methods [36, 28, 38, 17] struggle with noisy data, while multi-view image-based approaches leverage NeRF [18] and 3DGS [9]. Among these, NEF [35] reconstructs 3D edge fields from 2D maps using NeRF, followed by a fitting-then-clustering strategy to extract parametric edges. EMAP [12] learns an unbiased UDF for 3D edge modeling, addressing NEF’s inaccuracies, and employs a cluster-then-fitting strategy based on UDF gradients. More recently, EdgeGaussians [5] leverages 3DGS for efficient edge field reconstruction. Despite these advances, these works are constrained by a two-stage optimization paradigm, where the second stage of parametric edge extraction is prone to inaccuracies in edge representation and often produces fragmented and redundant edges. To address these issues, this paper proposes a one-stage parametric curve optimization pipeline, enabling more accurate and compact parametric edge extraction.
@@ 3 Method
Most existing methods [35, 12, 5] for reconstructing parametric 3D curves from multi-view edge images usually follow a two-stage approach: reconstructing a 3D edge representation (e.g., using NeRF, UDF or 3DGS) to get edge point clouds, then fitting parametric curves. While this approach is conceptually simple, it suffers from critical limitations. The two-stage paradigm leads to error accumulations. Since curve fitting heavily depends on the accuracy of the point cloud reconstruction, any noises from the first stage significantly degrade the final parametric curves. These issues also lead to lower efficiency. To address these limitations, we propose a one-stage optimization approach that directly optimizes parametric curves by the proposed curve-aware Gaussian splatting as illustrated in Fig. 2.
@@ 3.1 Problem Definition
Parametric Curves
In the framework, we represent each curve $\mathcal{C}_{j}$ as a parametric curve $\mathbf{c}_{j}(t)$, where $t\in[0,1]$ parameterizes the curve from start point ($t=0$) to end point ($t=1$). We use two types of parametric curves: cubic and first-order Bézier curves, while they are sufficient to describe most shapes encountered in most scenarios.
@@ 3.2 Curve-Aware Gaussian Splatting
The key challenge lies in designing a differentiable rendering function $f(\cdot)$. To achieve this, we propose a hybrid representation combining parametric curves with Gaussian primitives. We establish a bi-directional coupling between parametric curves and Gaussians, where each curve $c_{j}$ generates $N$ ($N$=12) Gaussians through uniform sampling as in Fig. 2. Gaussian attributes are determined by curve geometry:
@@
  • Mask: To address redundancy in overlapping regions and potential over-coverage in non-edge areas, we introduce a learnable importance mask $m^{j,i}\in[0,1]$ for each Gaussian following [11]. The mask operates as a gating mechanism during optimization, which means $\mathcal{G}_{j}^{i}$ is effectively suppressed when $m^{j,i}$ is close to 0. This attribute automatically indicates redundant or misplaced Gaussians along curve segments by gradient-driven mask decay.
As the result, the rendering function $f(\cdot)$ in Equ. 1 can be implemented by curve-aware Gaussian splatting. The splatting follows the standard 3D Gaussian Splatting (3DGS) pipeline but removes the color attribute, as the edge map is single-channel. Instead, alpha-weighted values are directly used to compute the rendered image and the rendering loss.
@@ 3.3 Adaptive Control of Curve-Aware Gaussians
While lacking prior knowledge of the number or locations of curves, we initialize the process with random Bézier parameterized curves. This approach leads to an optimization process that differs significantly from the original 3DGS. Specifically, the number of Gaussian primitives evolves dynamically, starting from a large set and gradually reducing to a smaller, more refined set during optimization. This behavior necessitates an adaptive strategy to ensure stable and accurate results. Four adaptive strategies are employed, as illustrated in the bottom part of Fig. 2, with details at below:
@@
During training, cubic Bézier curves are converted into first-order Bézier curves (line segments) if they are close to straight. Specifically, by uniformly sampling cubic Bézier curve $\mathbf{c}_{j}(t)$, we obtain a set of points $\{\mathbf{p}_{i}\}_{i=1}^{N}$, and then fit a straight line. The curve $\mathbf{c}_{j}(t)$ is replaced by a line segment $\mathbf{c}^{\prime}_{j}(t)$ if satisfying :
@@
where $\tau_{l}$ is a predefined error threshold, $\mathbf{P}_{j}^{0}$ and $\mathbf{P}_{j}^{3}$ are the start and end points of the cubic Bézier curve $\mathbf{c}_{j}(t)$, and $t\in[0,1]$. This transformation reduces computational complexity while preserving the geometric representation of curves.
@@
  • Line Merging: For two line segments $\mathbf{c}_{1}(t)$ and $\mathbf{c}_{2}(t)$, the angle $\theta$ between their directions and the spatial distance $d$ between their closest endpoints are computed. The lines are merged if $\theta<\tau_{la}\text{ and }d<\tau_{ld}$, where $\tau_{la}$ and $\tau_{ld}$ are predefined angular and spatial thresholds.
  • Cubic Bézier Curve Merging: For two adjacent Bézier curves $\mathbf{c}_{1}(t)$ and $\mathbf{c}_{2}(t)$, a new Bézier curve $\mathbf{c}^{\prime}(t)$ is fitted through a set of sampled points $\{\mathbf{p}_{i}\}_{i=1}^{N}$ from both curves. The two curves are merged and replaced with $\mathbf{c}^{\prime}(t)$ if $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{c}^{\prime}(t_{i})-\mathbf{p}_{i}\|_{2}<\tau_{b}$, where $\tau_{b}$ is a predefined error threshold.
This strategy not only simplifies the representation of the curves but also ensures they are compact, maintaining geometric accuracy and making it highly suitable for efficient curve-aware Gaussian splatting.
@@
A cubic Bézier curve $\mathbf{c}_{j}(t)$, with its control points $\mathbf{P}_{j,0}^{0},\mathbf{P}_{j,0}^{1},\mathbf{P}_{j,0}^{2},\mathbf{P}_{j,0}^{3}$, can be split at $t=s$ into two sub-curves using de Casteljau’s algorithm:
@@
  • Low mask detection: If the Gaussian with the smallest mask on a curve satisfies $m_{j}^{i}<\tau_{m}$, the curve is split at $t_{i-1}$ to retain $c_{1}$ and at $t_{i+1}$ to retain $c_{2}$, and the Gaussian point is removed.
@@
A curve $c_{j}(t)$ is dynamically removed if $o_{j}<\tau_{d}$, where $o_{j}$ denotes the opacity of the $j$-th curve, and $\tau_{d}$ is a predefined threshold. If the masks of all Gaussians on the curve are smaller than $\tau_{m}$, the curve will also be removed. This strategy ensures the removal of curves that are essentially transparent, thereby improving the overall quality of the representation.
@@ 3.4 Optimizing Curve-Aware Gaussians
where $\mathcal{L}_{1}$ and $\mathcal{L}_{D\text{-}SSIM}$ are supervised by the ground-truth image. However, this formulation suffers from gradient collapse in edge map optimization due to the sparse distribution of edge pixels, often leading to the transparency of all Gaussians converging to zero. To address this issue, we propose a novel edge-aware loss function $\mathcal{L}_{edge}$ that replaces $\mathcal{L}_{1}$ and balances the influence of edge and non-edge pixels in edge maps. The loss function is defined as:
$\mathcal{L}_{edge}=\frac{|M_{I}|}{|E_{I}|}\sum_{i\in N_{I}}\|I_{i}-\hat{I}_{i}\|_{2}^{2}+\frac{|N_{I}|}{|E_{I}|}\sum_{i\in M_{I}}\|I_{i}-\hat{I}_{i}\|_{2}^{2},$ (8)
where:
@@
  • $I_{i}$ and $\hat{I}_{i}$ are the pixel values in the rendered Gaussian image and the ground-truth edge map, respectively.
@@
$\mathcal{L}_{conn}=\sum_{i=1}^{M}\sum_{j=1}^{M}\mathbb{I}(\|P_{i}-P_{j}\|_{2}<\tau)\cdot\|P_{i}-P_{j}\|_{2}^{2},$ (9)
where $M$ is the total number of curves, $\tau$ is a distance threshold, and $\mathbb{I}(\cdot)$ is the indicator function that equals 1 if $\|P_{i}-P_{j}\|_{2}<\tau$ and 0 otherwise.
@@
$\mathcal{L}_{smo}=\sum_{j=1}^{M}\sum_{i=1}^{N_{j}-1}\|\mathbf{v}_{1}^{j,i}-\mathbf{v}_{1}^{j,i+1}\|_{2}^{2},$ (10)

### v3

@@ 1 Introduction
Figure 1: Our method achieves clean and accurate parametric curve reconstruction through a one-stage optimization utilizing the proposed curve-aware Gaussians. It eliminates the limitations of existing two-stage approaches, where the accuracy of the stage-2 curve fitting is heavily dependent on the quality of the initial point cloud reconstruction in stage-1.
@@
Parametric curves, as fundamental geometric primitives, play a vital role in structured 3D representations and geometric understanding [22, 38, 10, 17]. Their ability to compress geometric information while maintaining CAD compatibility [41] makes them indispensable for industrial applications. A critical challenge lies in reconstructing high-precision parametric curves from unstructured data, such as noisy point clouds or multi-view images, as sensor imperfections and incomplete observations [39, 37] inevitably degrade reconstruction quality.
@@
Early methods [53, 40, 56] for 3D curve reconstruction typically require clean 3D point clouds as input for subsequent fitting procedures. The recent NEF framework [52] first achieves 3D curve reconstruction directly from 2D images through a two-stage “edge point cloud reconstruction + parametric curve fitting” pipeline that combines neural radiance fields (NeRFs) [25] with RANSAC-based optimization [8]. However, this approach accumulates errors from the disconnected two-stage optimization process, making it sensitive to the noise of the reconstructed point cloud. EMAP [19] improves upon NEF by enhancing edge localization through the unsigned distance function (UDF) modeling and unbiased equation rendering. Though its “cluster-then-fitting” strategy reduces error propagation, the method remains fundamentally constrained by the point cloud quality given in the first stage. Notably, these approaches suffer from low computational efficiency due to NeRF optimizations [25]. While EdgeGaussians [6] accelerates reconstruction using Gaussian splatting [15], it still fails to overcome the inherent limitations of a two-stage fashion pipeline.
@@
The two-stage paradigm introduces two fundamental limitations. First, the parametric fitting process is strongly coupled with observation noises/errors, propagated from 2D edge detection to 3D point clouds through multi-view triangulation. It results in redundant branches or breaks in final curves (as shown in Fig. 1). Secondly, greedy iterative fitting (e.g., RANSAC [8]) produces local optima accumulation, causing curve redundancy to grow exponentially with scene complexity, which severely impacts downstream application efficiency. Eliminating the optimization gap between stages emerges as the key to robust reconstruction.
@@
While 2D edge maps are much easier to acquire, bypassing the explicit 3D edge point cloud reconstruction is a promising way to solve this problem. Direct one-stage curve optimization confronts two fundamental challenges: parametric representations inherently lack differentiable rendering capability, while neural rendering frameworks (NeRF/Gaussians) cannot preserve geometric continuity when approximating curves. We present a hybrid neural representation, CurveGaussian, that combines parametric curves with edge-oriented Gaussian components. By establishing bi-directional coupling, each Bézier curve segment precisely controls the spatial distribution of its associated Gaussian components. It enables differentiable rendering through splatting while significantly reducing the parameter count required for the optimization.
@@
Building upon this hybrid representation, this paper proposes a one-stage method for parametric curve reconstruction from multi-view edge maps. The core concept bypasses intermediate point cloud representations to directly perform differentiable optimization on the control points of Bézier curves, where each curve comprises multiple edge-oriented Gaussians. Specifically, each Gaussian component is bound to a specific edge segment of its corresponding curve, with positional and orientational attributes analytically derived from curve parameters. Rendering losses are backpropagated to update curve control points, enabling simultaneous exploitation of neural rendering’s photometric consistency constraints and parametric curves’ intrinsic geometric regularization. To further optimize curve topology during training for cleaner and more compact reconstructions, we propose a dynamic topology optimization strategy that adaptively refines edge structures through merging, splitting, and pruning operations. Additionally, we improve the optimization loss to prioritize the structural compactness of generated curves. Experiments demonstrate that our one-stage method increases 14.5% in reconstruction accuracy, improves computational efficiency by 33%, and reduces generated curve counts by 70.5% compared with two-stage SOTA alternatives, attaining more accurate and compact reconstructions. In summary, the contributions include:
@@ 2 Related Works
Early methods [33, 31, 5] reconstruct 3D line segments using Structure-from-Motion (SfM), involving line detection [26, 27, 45], matching [1, 28], and triangulation [3, 2, 5, 22]. Other works address line mapping by integrating graph clustering [13] and planar constraints [42] into SLAM systems [21, 23, 32]. NEAT [46] optimizes 3D line junctions using VolSDF [49] but is limited to straight lines and textured objects. Recent advances in 3D parametric curve reconstruction focus on two paradigms: point cloud-based and image-based approaches. Point cloud-based techniques [53, 40, 56, 24] extract parametric curves directly from edge point clouds, but their performance tends to degrade with noisy input. Alternatively, image-based methods have emerged as a promising solution, leveraging powerful radiance fields such as NeRF [25] and 3DGS [15]. These approaches benefit from the remarkable success of neural representations across various domains [12, 9, 55, 11], enabling high-quality 3D edge reconstruction from multi-view images that are typically pre-processed using edge detection [4, 36, 51]. Among these, NEF [52] and EMAP [19] reconstruct 3D edge fields from 2D images using the implicit radiance field, using clustering [8, 47, 48] to extract parametric edges. More recently, EdgeGaussians [6] leverages 3DGS for efficient edge field reconstruction. Despite these advances, these works are constrained by a two-stage optimization paradigm, where the second stage of parametric edge extraction is prone to inaccuracies in edge representation and often produces fragmented and redundant edges. To address these issues, this paper proposes a one-stage parametric curve optimization pipeline, enabling more accurate and compact parametric edge extraction.
@@ 3 Method
Most existing methods [52, 19, 6] for reconstructing parametric 3D curves from multi-view edge images usually follow a two-stage approach: reconstructing a 3D edge representation (e.g., using NeRF or 3DGS) to get edge point clouds, then fitting parametric curves. While this approach is conceptually simple, it suffers from critical limitations. The two-stage paradigm leads to error accumulations. Since curve fitting heavily depends on the accuracy of the point cloud reconstruction, any noise from the first stage significantly degrades the final parametric curves. These issues also lead to lower efficiency. To address these limitations, we propose a one-stage optimization approach that directly optimizes parametric curves by the proposed curve-aware Gaussian splatting as illustrated in Fig. 2.
@@ 3.1 Problem Definition
In the framework, we represent each curve $\mathcal{C}_{j}$ as a parametric curve $\mathbf{c}_{j}(t)$, where $t\in[0,1]$ parameterizes the curve from start point ($t=0$) to end point ($t=1$). We use two types of parametric curves: cubic and first-order Bézier curves, which are sufficient to describe most shapes encountered in most scenarios.
@@ 3.2 Curve-Aware Gaussian Splatting
The key challenge lies in designing a differentiable rendering function $f(\cdot)$. To achieve this, we propose a hybrid representation combining parametric curves with Gaussian primitives. We establish a bi-directional coupling between parametric curves and 3D Gaussians, where each curve $c_{j}$ generates $N$ ($N$=12) Gaussians through uniform sampling as in Fig. 2. Gaussian attributes are determined by curve geometry:
@@
  • Mask: To address redundancy in overlapping regions and potential over-coverage in non-edge areas, we introduce a learnable importance mask $m^{j,i}\in[0,1]$ for each Gaussian following [18]. The mask operates as a gating mechanism during optimization, which means $\mathcal{G}_{j}^{i}$ is effectively suppressed when $m^{j,i}$ is close to 0. This attribute automatically indicates redundant curve segments by gradient-driven mask decay.
As a result, the rendering function $f(\cdot)$ in Equ. 1 can be implemented by curve-aware Gaussian splatting. The splatting follows the standard 3DGS pipeline [15] but removes the color attribute, as the edge map is single-channel. Instead, alpha-weighted values are directly used to compute the rendered image and the rendering loss.
@@ 3.3 Adaptive Control of Curve-Aware Gaussians
While lacking prior knowledge of the number or locations of curves, we initialize the process with random Bézier parameterized curves. This approach leads to an optimization process that differs significantly from the original 3DGS. Specifically, the number of Gaussian primitives evolves dynamically, starting from a large set and gradually reducing to a smaller, more refined set during optimization. This behavior necessitates an adaptive strategy to ensure stable and accurate results. Four adaptive strategies are employed, as illustrated in the bottom part of Fig. 2, with details below:
@@
During training, cubic Bézier curves are converted into first-order Bézier curves (line segments) if they are close to straight. Specifically, by uniformly sampling the cubic Bézier curve $\mathbf{c}_{j}(t)$, we obtain a set of points $\{\mathbf{p}_{i}\}_{i=1}^{N}$, and then fit a straight line. The curve $\mathbf{c}_{j}(t)$ is replaced by a line segment $\mathbf{c}^{\prime}_{j}(t)$ if satisfying :
@@
where $\tau_{l}$ is a predefined error threshold, $\mathbf{P}_{j}^{0}$ and $\mathbf{P}_{j}^{3}$ are the start and end points of the cubic Bézier curve $\mathbf{c}_{j}(t)$, and $t\in[0,1]$. This transformation reduces computational complexity while preserving the precise geometry of the structure.
@@
  • Line Merging: For two line segments $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$, the angle $\theta$ between their directions and the spatial distance $d$ between their closest endpoints are computed. The lines are merged if $\theta<\tau_{la}\text{ and }d<\tau_{ld}$, where $\tau_{la}$ and $\tau_{ld}$ are predefined angular and spatial thresholds.
  • Cubic Bézier Curve Merging: For two adjacent Bézier curves $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$, a new Bézier curve $\mathbf{c}^{\prime}$ is fitted through a set of sampled points $\{\mathbf{p}_{i}\}_{i=1}^{N}$ from both curves. The two curves are merged and replaced with $\mathbf{c}^{\prime}$ if $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{c}^{\prime}(t_{i})-\mathbf{p}_{i}\|_{2}<\tau_{b}$, where $\tau_{b}$ is a predefined error threshold.
This strategy not only ensures the curves are compact but also maintains geometric accuracy, making it efficient for curve-aware Gaussian splatting.
@@
A cubic Bézier curve $\mathbf{c}_{j}$, with its control points $\mathbf{P}_{j,0}^{0},\mathbf{P}_{j,0}^{1},\mathbf{P}_{j,0}^{2},\mathbf{P}_{j,0}^{3}$, can be split at $t=s$ into two sub-curves using de Casteljau’s algorithm:
@@
  • Low mask detection: If the Gaussian with the smallest mask value on a curve satisfies $m_{j}^{i}<\tau_{m}$, indicating insufficient occupy, the curve is split at $t_{i-1}$ to retain $c_{1}$ and at $t_{i+1}$ to retain $c_{2}$, with the unreliable segment centered at $t_{i}$ removed.
@@
A curve $c_{j}$ is dynamically removed if $o_{j}<\tau_{d}$, where $o_{j}$ denotes the opacity of the $j$-th curve, and $\tau_{d}$ is a predefined threshold. If the masks of all Gaussians on the curve are smaller than $\tau_{m}$, the curve will also be removed. This strategy ensures the removal of curves that are essentially transparent, thereby improving the overall quality of the representation.
@@ 3.4 Optimizing Curve-Aware Gaussians
where $\mathcal{L}_{1}$ and $\mathcal{L}_{D\text{-}SSIM}$ are supervised by the ground-truth image. However, this formulation suffers from gradient collapse in edge map optimization due to the sparse distribution of edge pixels, often leading to the transparency of all Gaussians converging to zero.
To address this issue, we propose a novel edge-aware loss [52] function $\mathcal{L}_{edge}$ that replaces $\mathcal{L}_{1}$ and balances the influence of edge and non-edge pixels in edge maps. The loss function is defined as:
$\mathcal{L}_{edge}=\frac{|M_{I}|}{|E_{I}|}\sum_{i\in N_{I}}\|I_{i}-\hat{I}_{i}\|_{2}^{2}+\frac{|N_{I}|}{|E_{I}|}\sum_{i\in M_{I}}\|I_{i}-\hat{I}_{i}\|_{2}^{2}.$ (8)
@@
  • $I_{i}$ and $\hat{I}_{i}$ are the pixel values in the rendered image and the ground-truth edge map, respectively.
@@
$\mathcal{L}_{conn}=\sum_{i=1}^{C}\sum_{j=1}^{C}\mathbb{I}(\|P_{i}-P_{j}\|_{2}<\tau)\cdot\|P_{i}-P_{j}\|_{2}^{2},$ (9)
where $C$ is the total number of curves, $\tau$ is a distance threshold, and $\mathbb{I}(\cdot)$ is the indicator function that equals 1 if $\|P_{i}-P_{j}\|_{2}<\tau$ and 0 otherwise.
@@
$\mathcal{L}_{smo}=\sum_{j=1}^{C}\sum_{i=1}^{N_{j}-1}\|\mathbf{v}_{1}^{j,i}-\mathbf{v}_{1}^{j,i+1}\|_{2}^{2},$ (10)